OpenStudy (explainitlikeimfive):

What is the equation of y = x^3 with the given transformations? horizontal translation right 10 units and vertical translation down 12 units A. y = (x - 10)^3 - 12 B. y = (x + 10)^3 - 12 C. y = (x - 12)^3 + 10 D. y = (x + 12)^3 + 10

OpenStudy (queelius):

 So, when translating it 10 units to the right, that means where before the function y = x^3 was 0 at x = 0, now it should be 0 at x = 10. So, y = (x - 10)^3. Next, for every point on the graph, we want it to be shifted down 12 units. So, we can just substract 12 from it, y = (x - 10)^3 - 12.